OpenStudy (anonymous):

What is the answer to 3[(15-3)2\4]

OpenStudy (anonymous):

18?

OpenStudy (anonymous):

It's 18; i misread it.

OpenStudy (anonymous):

(15-3)2=24>24/4=6>3*6=18

OpenStudy (anonymous):

can you try the problem and show some work on what that was done